Find the slope of the line that goes through two points: (-5,8) and (-4,9)
Answer by danisanders021(4)

m=(y2-y2)/(x2-x1)
m=(9-(8))/(-4-(-5)*remember -(- equals +
therefore 9-(8)/(-4+5)
9-(8)/1 *anything over one equals itself like 2/1 equals 2
9-(8) =1 so 1/1 equals one
The answer is 1/1 or 1
y=1x+b